How do you find the mean and standard deviation for S2 - S1?

its asking for a general formula for to find the variance for two different time sets.
Answer #1

Mean (S1-S2) =E(S1-S2) =E(S1)-E(S2) =Mean (S1) -Mean (S2)

where